Problem 3: Calculate the current i(t). 250 700mH e 00000 to M 120cos (2001-897 404F ( Aco(2001-609 A
Answer #1

toomt 252 mun 4coscafoot-Go)A W 120cos (200t-8) +DE 120 cos (2007-8y ) w=200 radds 4 cos (2oot-60) A = 4 L-60A convert Llc vaApply KCl at v 120 L - - s = 456 1140 +25 -jia5 17=41-68 + Estas * 254570) 120 25+3140 (0 v [1-64xv141608=4-761-64.28 v= 1903